OpenStudy (anonymous):

Suppose that A is an nxn matrix with columns C1; C2; C3... Cn. If: 2C1 + 3C2 - C3 = 0 show that A is not invertible.

OpenStudy (helder_edwin):

are u allowed to use determinants?

OpenStudy (anonymous):

Yeah

OpenStudy (helder_edwin):

well then it is easy. the equation \[ \large 2C_1+3C_2-C_3=0 \] means that (at least) three columns are linearly dependant so the determinant is zero so the matrix is not invertible.

OpenStudy (anonymous):

Thanks

OpenStudy (helder_edwin):

u r welcome

OpenStudy (anonymous):

How would it be done without using determinants?

OpenStudy (helder_edwin):

the rank of a matrix is the dimension of both the row and column spaces. for a \(n\times n\) matrix to be invertible its rank has to be =n. since \[ \large 2C_1+3C_2-C_3=0 \] then the rank of this matrix is less than n. so the matrix cannot be invertible.